OpenStudy (anonymous):

Classify The Equation -3x=3-y 3+y+3x-2 Parallel Coincident Intersecting

OpenStudy (texaschic101):

-3x = 3 - y -3x - 3 = -y 3x + 3 = y can you please tell me the second equation..it is not clear

OpenStudy (anonymous):

-3x=3-y 3+y+3x-2

OpenStudy (texaschic101):

equation 2....there is no equal sign in there ?

OpenStudy (texaschic101):

3 + y + 3x - 2 = 0 y = -3x - 1 so we have : y = 3x + 3 and y = -3x - 1 different slopes, different y intercepts, INTERSECTING

OpenStudy (anonymous):

Thank You :) !